a body is resting on top of a hemispherical mound of ice of radius r if ice is frictionless then what minimum horizontal velocity must must be imparted to the body so that it leaves the mould without sliding over it
1
Expert's answer
2019-12-04T08:41:28-0500

Range = speed "\times" time

h "=0.5gt^2"

So,

"r=\frac{1}{2}gt^2\\or\\t=\sqrt\frac{2r}{g}"

Range = "u\times t=u\times \sqrt\frac{2r}{g}" "\ge r"


"u\ge \sqrt\frac{gr}{2}"